# Backdoor-based Explainable AI Benchmark for High Fidelity Evaluation of Attribution Methods

### v1

@@ 10 Proof
TABLE III: Comparison of accuracy among various ResNet-18 models Trojaned through Blend attack with triggers of varying visibility on CIFAR-10. The accuracy of the benign model (N/A) is also reported.
@@
TABLE IV: Comparison of accuracy among various ResNet-18 models Trojaned through Blend attack with triggers of varying visibility on GTSRB. The accuracy of the benign model (N/A) is also reported.
@@
TABLE V: Comparison of accuracy among various ResNet-34 models Trojaned through Blend attack with triggers of varying visibility on ImageNet. The accuracy of the benign model (N/A) is also reported.
@@
11 Experimental Setup
11.1 Trigger Patterns and Poisoned Samples
In our BackX benchmark, we incorporate both visible and invisible trigger patterns to provide a comprehensive evaluation, which allows us to thoroughly assess the capability of attribution methods in detecting both fixed visible patterns and input-specific invisible triggers. Figure 16 provides examples of the trigger patterns utilized in our experiments and the corresponding poisoned samples. In the Blend attack [24], we utilize fixed visible trigger patterns, consistent with Qi et al. [43]. In addition, we employ a pre-trained encoder network to generate input-specific invisible trigger patterns in the ISSBA attack [25]. The trigger patterns used in these attacks are depicted in the middle column of Figure 16. Given a set of input samples, as shown in the first column of Figure 16, the resulting poisoned samples are presented in the last column of Figure 16.
TABLE VI: Comparison of accuracy among various ResNet-18 models Trojaned through Blend attack with different poisoning rates on CIFAR-10.
@@ 11.1 Trigger Patterns and Poisoned Samples
TABLE VII: Comparison of accuracy among various ResNet-18 models Trojaned through Blend attack with different poisoning rates on GTSRB.
@@
TABLE VIII: Comparison of accuracy among various ResNet-18 models Trojaned through ISSBA attack on CIFAR-10 and GTSRB.
@@
TABLE IX: Accuracy comparison of ResNet-18, ResNet-32, ResNet-50 and ResNet-101 models Trojaned through Blend attack on ImageNet.
@@
Figure 17: The comparison of (a) attack success rate and (b) trigger recall across different model architectures using different attribution methods on ImageNet. A lower attack success rate indicates better results, while a higher trigger recall indicates better results.
11.2 Performance Comparison of Explained Models
In this part, we show the detailed performance of used models trained on different datasets. For all the Trojaned models under comparison, we select the target label as the first class for a single target backdoor attack on both the CIFAR-10 and ImageNet datasets, and the third class for the GTSRB dataset followed by Qi et al. [43].
Tables III and V display the accuracy comparison of Trojaned ResNet-18 models as the visibility of the Trojan trigger varies on the CIFAR-10 and GTSRB datasets. Similarly, Table V compares the accuracy of different Trojaned ResNet-32 models on ImageNet. In all cases, a consistent poisoning rate of $0.1$ was employed during model Trojaning. Notably, these tables illustrate that changes in trigger visibility do not lead to a marked decrease in accuracy on the clean dataset. Additionally, there is a minor reduction in accuracy on the poisoned dataset to ensure a fair comparison. These results provide compelling evidence that a sample containing the Trojan trigger can completely alter the model’s predictions, providing the ground truth of attributions. The first column in each table compares the accuracy of the benign model on the respective datasets, highlighting that Trojaned models only exhibit a slight performance compromise compared to their benign counterparts.
Table VII and Table VII present the accuracy changes in models Trojaned with different poisoning rates, ranging from 0.001 to 0.1, on the CIFAR-10 and GTSRB datasets. These tables demonstrate that models Trojaned with various poisoning rates only result in a subtle reduction in clean accuracy. It enables our experiments to maintain relatively consistent accuracy levels across different poisoning rates.
In Table IX, we show the accuracy comparison of ResNet models Trojaned through the ISSBA attack with a poisoning rate of $0.05$ on both CIFAR-10 and GTSRB datasets. We have ensured consistent standard accuracy and high accuracy on the poisoned dataset to enable a fair comparison. Furthermore, Table IX compares accuracy on both the poisoned and clean datasets across three different models.
The results highlight that Trojaned models exhibit the ability to adapt and fit the input distribution when optimizing the poisoned samples. As these Trojaned models converge during training on these samples, we are now equipped with the capacity to conduct a comprehensive and rigorous evaluation of attribution methods specifically within the context of Trojaned models. While other studies have demonstrated the presence of feature disparities in the hidden layers of Trojaned models within the latent space [43], our findings suggest that these models can still be considered reliable for evaluating explanations related to the model’s output layer. However, further exploring the impact of the disparity in evaluating attribution methods is left in our future work.
11.3 Hyperparameter Choice
@@
11.4 Experimental Software and Platform
@@
12 Experiments of Model Architectures
In Figure 17, we conduct a comparison of both attack success rate and trigger recall across different model architectures using various attribution methods. All three models were trained on ImageNet with a Trojan trigger of 0.5 visibility and a 0.1 poisoning rate. Table IX shows the detailed performance comparison of the three models. It can be observed that two CAM-based attribution methods and Grad exhibit sensitivity to changes in model architecture, indicating weaker stability. Conversely, most attribution methods show consistent performance across varying model architectures, demonstrating a stable behavior.
13 Extended Experiments of Overall Benchmarks
@@
13.2 Visual Inspection of Attributions

### v2

@@ A.7.1 Trigger Patterns and Poisoned Samples
Table 3: Comparison of accuracy among various ResNet-18 models Trojaned through Blend attack with triggers of varying visibility on CIFAR-10. The accuracy of the benign model (N/A) is also reported.
@@
Table 4: Comparison of accuracy among various ResNet-18 models Trojaned through Blend attack with triggers of varying visibility on GTSRB. The accuracy of the benign model (N/A) is also reported.
@@
Table 5: Comparison of accuracy among various ResNet-34 models Trojaned through Blend attack with triggers of varying visibility on ImageNet. The accuracy of the benign model (N/A) is also reported.
@@
A.7.2 Performance Comparison of Explained Models
In this part, we show the detailed performance of used models trained on different datasets. For all the Trojaned models under comparison, we select the target label as the first class for a single target backdoor attack on both the CIFAR-10 and ImageNet datasets, and the third class for the GTSRB dataset followed by Qi et al. (Qi et al., 2022).
Table 6: Comparison of accuracy among various ResNet-18 models Trojaned through Blend attack with different poisoning rates on CIFAR-10.
@@ A.7.2 Performance Comparison of Explained Models
Table 7: Comparison of accuracy among various ResNet-18 models Trojaned through Blend attack with different poisoning rates on GTSRB.
@@
Table 8: Comparison of accuracy among various ResNet-18 models Trojaned through ISSBA attack on CIFAR-10 and GTSRB.
@@
Table 9: Accuracy comparison of ResNet-18, ResNet-32, ResNet-50 and ResNet-101 models Trojaned through Blend attack on ImageNet.
@@
Tables 3 and 5 display the accuracy comparison of Trojaned ResNet-18 models as the visibility of the Trojan trigger varies on the CIFAR-10 and GTSRB datasets. Similarly, Table 5 compares the accuracy of different Trojaned ResNet-32 models on ImageNet. In all cases, a consistent poisoning rate of $0.1$ was employed during model Trojaning. Notably, these tables illustrate that changes in trigger visibility do not lead to a marked decrease in accuracy on the clean dataset. Additionally, there is a minor reduction in accuracy on the poisoned dataset to ensure a fair comparison. These results provide compelling evidence that a sample containing the Trojan trigger can completely alter the model’s predictions, providing the ground truth of attributions. The first column in each table compares the accuracy of the benign model on the respective datasets, highlighting that Trojaned models only exhibit a slight performance compromise compared to their benign counterparts.
Table 7 and Table 7 present the accuracy changes in models Trojaned with different poisoning rates, ranging from 0.001 to 0.1, on the CIFAR-10 and GTSRB datasets. These tables demonstrate that models Trojaned with various poisoning rates only result in a subtle reduction in clean accuracy. It enables our experiments to maintain relatively consistent accuracy levels across different poisoning rates.
In Table 9, we show the accuracy comparison of ResNet models Trojaned through the ISSBA attack with a poisoning rate of $0.05$ on both CIFAR-10 and GTSRB datasets. We have ensured consistent standard accuracy and high accuracy on the poisoned dataset to enable a fair comparison. Furthermore, Table 9 compares accuracy on both the poisoned and clean datasets across three different models.
The results highlight that Trojaned models exhibit the ability to adapt and fit the input distribution when optimizing the poisoned samples. As these Trojaned models converge during training on these samples, we are now equipped with the capacity to conduct a comprehensive and rigorous evaluation of attribution methods specifically within the context of Trojaned models. While other studies have demonstrated the presence of feature disparities in the hidden layers of Trojaned models within the latent space (Qi et al., 2022), our findings suggest that these models can still be considered reliable for evaluating explanations related to the model’s output layer. However, further exploring the impact of the disparity in evaluating attribution methods is left in our future work.
A.7.3 Hyperparameter Choice
@@
A.7.4 Experimental Software and Platform
@@
A.8 Re-Calibrated Attribution and Contrastive Output
Recent research has explored alternatives to enhance reliability of baseline attribution results (Wang & Wang, 2022), (Yang et al., 2023b). Here, we conduct comparative experiments for such choices. Re-calibrated attributions (Wang & Wang, 2022) have been introduced to enhance integration-based methods by estimating attributions from identified valid references. In Fig. 10, we compare the re-calibrated attributions and original attributions across integration-based attribution methods, including IG-SG, IG-Uni, AGI and LPI. It is observed that re-calibrated attributions further enhance the performance of the original attributions on CIFAR-10, GTSRB and ImageNet datasets. In addition, the performance of re-calibrated attributions is observed to degrade with decreasing mean input pixel values, aligning with our Observation 2. This demonstrates the sensitivity of integration-based methods to variations in the reference values. Despite the enhancement achieved through re-calibration, it remains unique to integration-based attribution methods. To maintain evaluation consistency across different methods, we retain the use of original attributions in the integration methods for the subsequent experiments.
Figure 11 offers a comparison between attributions computed by our established consistent setup and contrastive outputs (Wang & Wang, 2022) on CIFAR-10, GTSRB, and ImageNet. It is notable that explaining contrastive output yields similar performance compared to explaining manually selected outputs, particularly evident in cases involving FullGrad and integration-based methods. This observation underscores the superiority of contrastive output over mutual selection. However, explaining contrastive output in CAM and Gradient-based methods does not achieve comparable performance. The following observation is derived from the above analysis.
Observation 3.1 The re-calibration technique consistently improves the performance of integration-based attribution methods. Attribution derived from explaining contrastive model output can achieve competitive results, particularly for CAM-based and integration-based methods.
A.9 Extended Experiments of Output Choice
In this part, we provide more experimental results of attribution methods in explaining different objects.
[…]
A.10 Extended Experiments of Overall Benchmarks
@@
A.10.1 Results of Fractional Probability Change
A.10.2 Results of Other Attribution Methods
In this study, we primarily aim to enhance model reliability, which led us to focus on attribution methods with established robustness. Consequently, SHAP-based and perturbation-based approaches were excluded from the main benchmark, as previous works have highlighted their reliability limitations (Shrikumar et al., 2017; Sundararajan et al., 2017). Furthermore, perturbation-based methods are generally more computationally intensive compared to propagation-based alternatives. However, to assess their potential contributions, we extend our evaluation to include additional attribution methods that were not the primary focus of our study. These methods include LIME (Ribeiro et al., 2016), LRP (Binder et al., 2016), DeepLIFT (Shrikumar et al., 2017), KernelSHAP (Lundberg & Lee, 2017), GradientSHAP (Lundberg & Lee, 2017), HSIC (Novello et al., 2022), and SMDLShap (Chen et al., 2024), alongside IG, SG, and FG, using ImageNet on MobileNet-V2.
Figures 15(a) and 15(b) present a comparison of the attack success rate and trigger recall for both focused and unfocused attribution methods. The results indicate that the unfocused methods tend to perform less favorably than the focused methods. Additionally, we observe that SHAP-based methods exhibit high sensitivity to the choice of feature set, further complicating their application.
Figure 15: Comparative analysis of focused and unfocused attribution methods using (a) attack success rate and (b) trigger recall on MobileNetV2 trained on ImageNet. A lower attack success rate signifies better performance, while a higher trigger recall indicates better performance.
A.10.3 Extended Experiments on Model Architectures and Depths
In this part, we examine the impact of different model architectures and model depths on attribution performance.
Figure 16: The comparison of different attribution methods using (a) attack success rate and (b) trigger recall on different model architectures trained on ImageNet. A lower attack success rate indicates better results, and a higher trigger recall indicates better results.
Figure 17: The comparison of different attribution methods using (a) attack success rate and (b) trigger recall on ResNet of depth 18, 34, 50, and 101 trained on ImageNet. A lower attack success rate indicates better results, and a higher trigger recall indicates better results.
1) Different Model Architectures: Figure 16 compares the attack success rate and trigger recall across various model architectures, including MobileNetV2 (Sandler et al., 2018), VGG16 (Simonyan & Zisserman, 2015), and ResNet-50 (He et al., 2016), using different attribution methods. All models were trained on ImageNet with a Trojan trigger of 0.5 visibility and a 0.1 poisoning rate. The results show that two CAM-based attribution methods and input gradient attribution methods are sensitive to changes in model architecture, demonstrating lower stability. In contrast, other evaluated attribution methods exhibit consistent performance across different architectures, suggesting higher stability. Additionally, ResNet-50, when compared to MobileNetV2 and VGG16, proves to be more challenging to explain, as evidenced by its higher attack success rate and lower trigger recall.
2) Different Model Depth: Figure 17 compares attribution methods on ResNet models of varying depths, assessing both attack success rate and trigger recall on ImageNet. Similarly, the two CAM-based methods and input gradient attributions show instability when explaining models of different depths. On the other hand, the other attribution methods exhibit consistent performance across models of varying depths. Overall, model depth does not significantly complicate the task of attribution.
Based on these evaluations, we make the following observations:
Observation 5. CAM-based and vanilla input gradient attributions are sensitive to model changes. In contrast, most other attribution methods show consistent performance across different models. Overall, more complex architectures generally increase the difficulty of attribution, whereas changes in model depth pose less of a challenge.
A.11 Extended Benchmarking Results on Textual Analysis
Most attribution methods have been primarily focused on the image domain. In this work, we extend our benchmark to the textual domain, further demonstrating the applicability of BackX. In this section, we first outline the benchmarking setup and then provide a comprehensive evaluation of attribution methods in the context of textual analysis.
A.11.1 Textual Benchmarking Process
To benchmark attribution methods in the textual domain, we construct Trojaned language models using the BadNet attack (Gu et al., 2019), as implemented by OpenBackdoor (Cui et al., 2022). We consider four different trigger patterns–”cf”, ”mn”, ”bb”, and ”tq”-which are randomly injected at predefined locations within the input text while modifying the corresponding training labels to generate poisoned samples. We fine-tune a BERT-based model on three poisoned sentiment analysis datasets: SST-2 (Socher et al., 2013), IMDB (Maas et al., 2011), and Amazon Reviews (He & McAuley, 2016). Specifically, we employ a pre-trained BERT-based model (Kenton & Toutanova, 2019) with 12 transformer layers, 768 hidden units, and 12 attention heads as the victim model for textual Trojaning. Table 10 reports the performance of the Trojaned BERT models across the three datasets. The results indicate that neural Trojans cause significant alterations in the model’s predictions, thereby confirming the fidelity of the subsequent benchmarking analysis.
For the benchmarked attribution methods, we exclude class activation maps (CAM) typically used in image classification tasks to highlight discriminative regions, as such maps are not directly applicable to textual classification. Instead, we focus on gradient-based and integral-based attribution methods, which are better suited for generating attributions in the textual domain. Unlike image-based models, where the input data is manipulated at the pixel level, textual inputs are first tokenized into discrete token IDs. This tokenization results in a representation that is not differentiable, as the input IDs are discrete and do not support gradient-based manipulation. Consequently, direct attribution to the raw input text is not feasible. To overcome this limitation, we compute attributions for the input features after the embedding layer, where the token IDs are transformed into continuous and differentiable representations. Thus, in our BackX benchmarking, we estimate attributions based on the embeddings of the input text.
Table 10: Accuracy comparison of Bert-based models Trojaned through BadNet attack on SST-2, IMDB and Amazon Review datasets. The accuracy on both poisoned and clean testing set are presented.
Test set Sentiment Analysis Datasets
SST-2 IMDB Amazon
Poisoned $0.924$ $0.935$ $0.961$
Clean $1.000$ $0.938$ $0.982$
Figure 18: Benchmarking of attribution methods under BadNet attack using the attack success rate metric on Bert-based models trained on SST-2, IMDB and Amazon datasets. Lower curves indicate better results.
Figure 19: Benchmarking of attribution methods under BadNet attack using the trigger recall metric. Higher curves indicate better results.
A.11.2 Evaluation on Textual Analysis
[…]
The results indicate that Integrated Gradients (IG), InputGrad (Grad), and SmoothGrad (SG) exhibit the most consistent and robust performance across all three datasets. Notably, IG’s performance in the textual domain differs significantly from that in the image domain, underscoring its high reliability for generating explanations. This enhancement in reliability is attributed to the more reliable selection of reference inputs by IG. Specifically, in contrast to the vision domain, IG’s use of zero-valued vectors as reference inputs accurately captures the absence of features in the textual domain. This distinction is reflected in the performance gap between IG and its variants, which employ alternative reference inputs such as Gaussian noise, uniform noise, and squared Gaussian noise. Moreover, due to the uncertainty introduced by perturbations transferring from the input to the feature space, IG’s advantage becomes more pronounced when attributing features in the embedding layer, leading to superior performance.
Similarly, Grad demonstrates competitive results compared to SG, showing substantial improvements in the textual domain. SG’s reduced effectiveness in text attribution can be attributed to its reliance on noisy gradients, which struggle to capture complex interactions in feature space. Due to the lower perturbation amplitude, its performance remains close to that of Grad. These findings suggest that perturbation-based attribution methods fail to enhance explanations in the textual domain effectively. Moreover, the observed variability in attribution reliability across domains underscores the necessity of tailoring attribution methods to improve explanation faithfulness. The benchmarking results on textual analysis further highlight the potential for extending our framework beyond textual data to other modalities, such as audio, time-series, and graphs.
Figure 20: Attribution visualization for explaining a text sample from the SST-2 Dataset. Given an input sample, the poisoned sample using trigger “mn” alters the sentiment prediction from negative to positive. The importance of each input word, as estimated by different attribution methods, is visualized. Brighter colors indicate more important words.
In Fig. 20, we present visualization results of the attributions used to explain sentiment prediction outcomes on SST-2 dataset. For a given input sample, the poisoned sample containing the ”mn” trigger causes a shift in the sentiment prediction from negative to positive. IG and Grad effectively attribute this prediction shift to the trigger pattern, whereas other attribution methods exhibit compromised performance in explaining this shift.
Observation 6. For textual analysis, IG and Grad achieve superior performance. IG, leveraging a zero reference, effectively mitigates challenges associated with reference selection by accurately capturing the absence of features. In contrast, perturbation-based attribution methods suffer degradation due to uncertainty in the feature space. The observed variability in attribution reliability across domains highlights the need for tailored attribution methods that account for domain-specific characteristics.
Figure 21: Comparative analysis of defending neural backdoors using different attribution methods on MobileNet-V2 model trained on (a) CFAIR-10 and (b) GTSRB datasets. We using attributions. For each poisoned sample, pixels corresponding to the top 50% highest attribution scores are restored to their clean state. Lower attack success rate indicates better defense performance.
A.12 Extended Experiments of Backdoor Defense
A.12.1 Explainability under Diverse Attack Methods
We investigate the effectiveness of attribution methods in identifying and mitigating various types of neural Trojans. Beyond the Blend and ISSBA attacks, we extend our analysis to additional backdoor methods, including SIG (Barni et al., 2019), WaNet (Nguyen & Tran, 2021), and Adap-Blend (Qi et al., 2022). SIG utilizes sinusoidal signals as triggers, seamlessly blending into textures to evade spatial filtering defenses. WaNet introduces geometric warping rather than pixel-level modifications, rendering backdoors visually imperceptible. Adap-Blend enhances stealth by reducing feature separability in the latent space, making Trojaned features less distinguishable from normal ones.
Figure 21 presents a comparative evaluation of attribution methods on MobileNet-V2 trained on CIFAR-10 and GTSRB. In this evaluation, we recover 50% of the poisoned sample pixels based on the attribution scores, restoring these pixels to their clean-state counterparts from unaltered samples. The results indicate that CAM-based attribution methods are more effective in detecting triggers with fixed and consistent spatial patterns, as evidenced by their superior performance on Blend, Adap-Blend, and WaNet. In contrast, gradient-based and integral-based attribution methods excel in identifying discrete and imperceptible triggers, such as those used in SIG and ISSBA attacks.
However, attribution methods struggle to effectively neutralize the Adap-Blend backdoor. This limitation arises because Adap-Blend utilizes image-wide trigger patterns and applies regularization that encourages individual features affected by the trigger to independently influence the model’s predictions. Consequently, even with a recovery rate of 50%, Adap-Blend maintains a relatively high attack success rate. These findings highlight the necessity of designing tailored attribution methods that account for the distinct characteristics of different backdoor strategies.
Observation 7. CAM-based methods excel at detecting fixed spatial triggers, while gradient- and integral-based methods are more effective against discrete and imperceptible ones, highlighting the need for tailored attribution techniques for backdoor defending.
@@ A.12.1 Explainability under Diverse Attack Methods
A.12.2 Results of Attack Success Rate
A.13 Discussion and Limitation
In this section, we discuss the position of BackX in attribution evaluation and clarify its contribution to the fidelity benchmark alongside limitations.
What number and diversity of Trojans are sufficient for BackX to deliver confident benchmarking results of attributions? Similar to all existing attribution benchmarks, our underlying goal is to provide a necessary guarantee for rejecting unreliable attribution methods. In this context, any number of Trojans can provide sufficient confidence in evaluating attribution reliability. From the perspective of necessary conditions, since the fundamental requirement of any explanation method is to remain faithful to the model’s prediction, all Trojans introduced in our framework are sufficient to provide strong evidence against unreliable attribution methods. If an attribution method fails to provide faithful explanations under any controlled Trojaning scenario, it should be rejected from real-world deployment. Compared to prior benchmarks (e.g., CLEVR-XAI, Null Feature, and DiPart), our method offers a fully controllable setting to test explanation reliability and further strengthens this ability by providing a tighter theoretical guarantee, thus offering a more solid reason for rejecting unreliable attribution methods, especially in high-stakes applications. From the perspective of sufficient conditions, while arguably no benchmark can guarantee that an attribution method is universally reliable, we strive to ascertain the sufficiency of BackX across diverse domains by extending beyond simple patch-based triggers with diverse trigger patterns, multiple fidelity metrics, and different modalities, thereby benchmarking attribution methods under varied and increasingly challenging conditions. This progressive expansion helps toward the sufficiency of our evaluation, while we acknowledge that sufficiency remains inherently limited across all existing benchmarks and explicitly aim to push this boundary further. In summary, our method offers a stronger reliability guarantee than existing benchmarks by enforcing a necessary condition for attribution methods. If an attribution method fails under our controlled setting, it provides a clear reason to reject its reliability. This necessity guarantee represents a central advantage of our approach over prior works.
Does the mechanism of neural Trojans provide attribution fidelity? Unlike prior works that detect backdoor triggers using attribution methods (Lin et al., 2021; Li et al., 2023), BackX is a high-fidelity benchmark established within rigorous fidelity criteria through tailored strategies with tighter theoretical guarantees. Therefore, it is not merely the existence of neural Trojans, but BackX’s designs enable Trojans to serve as a pathway for achieving a reliable high-fidelity attribution benchmarking. Specifically, our framework formalizes fidelity as verifiable criteria and provides stronger guarantees by mitigating leakage and enforcing restoration so that prediction shifts are causally attributable. With such fidelity guarantees, BackX introduces a standardized evaluation setup across attribution paradigms, reducing confounding factors from post-processing and output forms. It also extends coverage substantially, testing fifteen attribution methods under five Trojan families across both vision and text. This comprehensive setup allows us to draw broader and more reliable conclusions. In Table 11, we report pairwise Kendall’s $\tau$ between benchmark rankings for the seven attribution methods SG, AGI, IG, IG-SG, FG, IxG, and GCAM. The results show that BackX is broadly consistent with DiFull, DiffID, and ROAR, whereas NeuronTrojans aligns weakly with BackX and also exhibits reduced consistency with the other benchmarks. The results provide compelling evidence that a backdoored model on its own does not deliver benchmarking fidelity. Fidelity arises only when backdoors are embedded within a framework that enforces formal criteria, theoretically grounded procedures, and a standardized evaluation protocol. BackX treats backdoors as instruments to isolate attribution effects and to satisfy fidelity requirements, rather than as mere objects of detection.
Table 11: Pairwise Kendall’s $\tau$ between benchmark rankings on seven attribution methods. Higher values indicate stronger agreement in method rankings.
Benchmarks Kendall’s $\tau$
BackX vs DiFull 0.333
BackX vs DiffID 0.619
BackX vs DiffROAR 0.714
BackX vs NeuronTrojans 0.048
NeuronTrojans vs DiFull -0.238
NeuronTrojans vs DiffID 0.429
NeuronTrojans vs DiffROAR 0.143
Limitation. While BackX evaluates attribution methods under a diverse set of trigger patterns, it does not explicitly measure the stability of their performance across these variations. Moreover, although BackX offers a rigorous and necessary test for rejecting unreliable attribution methods, strong performance under its controlled settings may not necessarily imply full reliability in complex, real-world scenarios. Expanding BackX to better capture complex real-world attribution patterns and to generalize across broader domains remains an important direction for future research.
A.14 Visual Inspection of Attributions
In Figure 24, we provide visualizations of employing absolute and original attributions for different methods. There is a notable disparity in the highlighted regions of input samples between the two post-processing techniques. This underscores the susceptibility of visual outcomes to ambiguity, emphasizing the necessity of quantitative assessments to provide clarity and context. This problem is more pronounced for smaller input images due to the limited number of pixels. Although large ImageNet size images may yield more consistent attribution maps, it is crucial to recognize that their relative feature importances undergo shifts, often reflected in attributions alternating between negative and positive values. This begs for an enquiry into the right choice between the original and absolute values for benchmarking.